A binary code (leading zeros allowed) consisting of 6 digits is randomly generated by a computer. What is the probability that a code is generated with an alternating pattern of 00 and 11?

Write your answer as a fraction 1a1a and enter only the denominator aa below.
Answer #1

here total number of  6 digits variables =2^6 =64

number of ways 00 and 11 can alteranate =2

therefore  probability that a code is generated with an alternating pattern of 00 and 11 =2/64 =1/32